# Discovering Software Parallelization Points Using Deep Neural Networks

### v1

@@ 3.4 Statistical Summary of Model Performance
A detailed statistical evaluation was conducted to assess the average behavior and consistency of both DNN and CNN models under different levels of data compression, specifically with 95%, 90%, 85%, and 80% of retained PCA variance. Tables 3 to 6 present key performance metrics, including mean test accuracy, standard deviation, median, best and worst outcomes across 30 executions, and 95% confidence intervals calculated using Student’s t-test.
Tables 3 and 4 summarize the mean, standard deviation, and median of the test accuracy for each model across all training proportions. The DNN achieved its highest average accuracy (94.04%) with the 85% configuration, which also presented the lowest standard deviation (1.66%) and a median of 94.16%, indicating stable and consistent performance. In contrast, the CNN achieved its peak average accuracy of 93.09% with the 90% configuration and a median of 94.25%, although with slightly greater variability across runs.
@@ 3.5.1 DNN Best-Case Performance (PCA 85%)
Figure 6 presents the training behavior of the best DNN model. It contains two subplots: one showing training and validation accuracy, and the other showing training and validation loss across epochs.
@@
The model reached a final training accuracy of 88.61% and a validation accuracy of 95.00%, indicating effective generalization. The smooth convergence of both curves suggests stable training without overfitting.
Table 7 shows the corresponding confusion matrix, revealing balanced performance across both classes, with a low number of misclassifications.
@@
The classification report for this execution is shown in Table 8. The model achieved high precision, recall, and F1-scores for both classes, with macro and weighted averages all at 0.97.
@@ 3.5.2 CNN Best-Case Performance (Full Features)
Figure 7 shows the training and validation performance of the best CNN model. Similar to the DNN, it includes two subplots: one displaying accuracy curves and the other showing loss curves. The CNN reached a training accuracy of 88.11% and a validation accuracy of 93.17%. The convergence of both curves suggests stable training without the presence of overfitting.
@@
The confusion matrix in Table 9 shows excellent classification performance, with zero false negatives for class 0 and a small number of false positives in class 1.
@@ 3.6 Worst-Case Evaluation of DNN and CNN Models
While the previous subsection highlighted the best-performing executions, this section focuses on the opposite end of the performance spectrum. Here, we examine the worst-performing execution for both the DNN and CNN models. Importantly, all models were trained under the same conditions; the variations in performance stem solely from differences in initialization and training dynamics. As in the previous analysis, the evaluation includes training plots, confusion matrices, and classification reports.
@@ 3.6.1 DNN Worst-Case Performance (Full Features)
Figure 8 shows the training behavior of the worst-performing DNN model, which used 100% of the original features.
@@
It contains two subplots: one for training and validation accuracy, and another for training and validation loss. While the model achieved a training accuracy of 88.50%, the validation accuracy dropped significantly to 62.67%, reflecting strong overfitting. The large gap between training and validation performance indicates that the model memorized the training data but failed to generalize effectively to unseen samples.
Table 11 presents the confusion matrix, revealing a strong class imbalance in prediction. The model failed to correctly classify most instances of class 0 (Undefined Loop), with 244 misclassifications.
@@
The classification report in Table 12 further emphasizes the imbalance, with the model achieving a perfect recall for class 1 but very poor performance on class 0. The overall accuracy was 59.33%.
@@ 3.6.2 CNN Worst-Case Performance (Full Features)
The classification report in Table 14 confirms the poor generalization of this model, particularly for class 1. While it reached high precision on that class, the recall and F1-score were extremely low, resulting in an overall accuracy of 55.83%.

### v2

@@ 3.4 Statistical Summary of Model Performance
A detailed statistical evaluation was conducted to assess the average behavior and consistency of both DNN and CNN models under different levels of data compression, specifically with 95%, 90%, 85%, and 80% of retained PCA variance. Tables 3 to 6 present key performance metrics, including mean test accuracy, standard deviation, median, best and worst outcomes across 30 executions, and 95% confidence intervals calculated using Student’s t-test, providing a comprehensive overview of each model’s performance variability and allowing a direct comparison of stability under varying amounts of retained information.
Tables 3 and 4 summarize the mean, standard deviation, and median of the test accuracy for each model across all training proportions. The DNN achieved its highest average accuracy (94.04%) with the 85% configuration, which also presented the lowest standard deviation (1.66%) and a median of 94.16%, indicating stable and consistent performance across runs and limited fluctuation between executions. Similarly, the CNN reached its peak average accuracy of 93.09% with the 90% configuration, accompanied by a median of 94.25%, although the variability across runs was slightly higher compared to the DNN, as reflected by its standard deviation. These results highlight the relative stability of both models, the slight differences in sensitivity to PCA variance retention, and provide a clear picture of how the proportion of retained features affects predictive consistency and overall performance.
@@ 3.5.1 DNN Best-Case Performance (PCA 85%)
Figure 6 presents the training behavior of the best DNN model. It contains two subplots: one showing training and validation accuracy, and the other showing training and validation loss across epochs. The model reached a final training accuracy of 88.61% and a validation accuracy of 95.00%, indicating effective generalization. The smooth convergence of both curves suggests stable training without overfitting.
@@
Table 7 shows the corresponding confusion matrix, revealing balanced performance across both classes. The model correctly classified most samples, with only a small number of misclassifications, 11 for class 0 and 8 for class 1, indicating strong overall predictive capability. The classification report for this execution is shown in Table 8. The model achieved high precision, recall, and F1-scores for both classes, demonstrating balanced performance and strong predictive capability. Both macro and weighted averages are 0.97, reflecting consistent accuracy across the dataset. The support values indicate that the evaluation included 323 samples of Undefined Loop (0) and 277 samples of Independent Loop (1), confirming that the model performed reliably across classes of differing sizes. Overall, these results highlight the DNN’s robustness in correctly classifying the majority of samples and maintaining high performance metrics throughout.
@@ 3.5.2 CNN Best-Case Performance (Full Features)
Figure 7 shows the training and validation performance of the best CNN model. Similar to the DNN, it includes two subplots: one displaying accuracy curves and the other showing loss curves. The CNN reached a training accuracy of 88.11% and a validation accuracy of 93.17%, indicating effective learning of the training data while generalizing well to unseen samples. The convergence of both curves and the relatively small gap between training and validation metrics suggest stable training, consistent improvement over epochs, and the absence of overfitting.
@@
The confusion matrix in Table 9 shows excellent classification performance, with zero false positives for class 0 and a small number of false negatives in class 1.
@@ 3.6 Worst-Case Evaluation of DNN and CNN Models
While the previous subsection highlighted the best-performing runs, this section focuses on the opposite end of the performance spectrum. Here, we examine the worst-performing runs for both the DNN and CNN models. As in the previous analysis, the evaluation includes training plots, confusion matrices, and classification reports, allowing for a detailed assessment of the models. By analyzing the worst-case performance, we can gain a deeper understanding of the limitations of each model and the scenarios in which they fail to generalize effectively.
@@ 3.6.1 DNN Worst-Case Performance (Full Features)
Figure 8 shows in detail the training behavior of the worst-performing DNN model, which was configured using 100% of the original features as input. The figure is divided into two subplots: the first illustrates the evolution of training and validation accuracy over the epochs, and the second presents the curves of training and validation loss. At the end of training, the model reached a relatively high training accuracy of 88.50%, but its validation accuracy dropped considerably to 62.67%. This sharp decline clearly reflects strong overfitting, since the model demonstrated good performance on the training set but failed to reproduce similar results on the validation set. The large gap between training and validation curves shows that the network tended to memorize the training data instead of learning patterns capable of generalizing to new, unseen samples. This discrepancy highlights the sensitivity of the DNN model to the specific characteristics of the training data and reinforces the importance of simultaneously analyzing both training and validation metrics when assessing the overall reliability and robustness of the model.
@@
Table 11 presents the confusion matrix of the worst-performing DNN model, revealing a strong class imbalance in prediction. The model failed to correctly classify most instances of class 0 (Undefined Loop), with 244 misclassifications, while class 1 (Independent Loop) was classified correctly in all cases. This imbalance is further highlighted in the classification report shown in Table 12, where the model achieved perfect recall for class 1 but very low recall for class 0, resulting in a low F1-score for that class and an overall accuracy of 59.33%.
@@ 3.6.2 CNN Worst-Case Performance (Full Features)
The classification report in Table 14 confirms the poor generalization of this model, particularly for class 1 (Independent Loop). While it reached a high precision of 0.93 for this class, the recall dropped drastically to 0.05, and the F1-score was only 0.09, indicating that the model failed to correctly identify most instances of this class. In contrast, class 0 (Undefined Loop) achieved perfect recall (1.00) but low precision (0.55), showing that many predictions were misclassified as this class. Overall, these imbalances resulted in a low overall accuracy of 55.83%, with macro and weighted averages also reflecting poor performance (0.40 and 0.42 F1-score, respectively). This highlights the model’s inability to handle class-specific distinctions effectively under worst-case conditions.